The front of a house is a triangle that is as wide as it is tall.It's area is 98ft^2.find the height and the base.Please help
Answer by mananth(16949)

Area of triangle = 1/2 * base * height
height & base are equal. let them be x
98=1/2 * x^2
196=x^2
14= x which is the height and base

ananth